# Thermodynamic Variational Principle Unifying Gravity and Heat Flow

### v1

@@ Abstract
Equilibrium systems, even under weak gravity, are governed by thermodynamic functions and associated variational principles, such as the minimum free energy principle, which uniquely select the stable state [8, 9, 10, 11]. However, the presence of a sustained heat flux establishes a NESS where standard equilibrium principles fail [7]. Developing a general predictive framework for NESS remains a central challenge, addressed by various approaches such as large deviation theory [12, 13] and stochastic thermodynamics [14, 15]. As a variational principle in weak NESS, the minimum entropy production principle, originating from linear irreversible thermodynamics [16, 17], was proposed for the selection of a steady current [18]. Despite the mathematically precise formulation of the minimum entropy production principle [19, 20], it is still challenging to predict the true steady state among multiple steady states.
@@
We consider $N$ particles of mass $m$ confined in a rectangular container of volume $V$, height $L$, and cross-sectional area $A=V/L$. See Fig. 1. The vertical coordinate $x$ runs from the bottom wall ($x=0$) to the top wall ($x=L$). Uniform gravity $g$ acts in the negative $x$ direction, resulting in a gravitational potential energy $mg(x-x_{\rm o})$ relative to an arbitrary reference point $x_{\rm o}$. The walls at 00 and $L$ are maintained at constant temperatures $T_{1}$ and $T_{2}$, respectively. We define the temperature difference parameter $\Xi\equiv T_{2}-T_{1}$. The total particle number $N$ and volume $V$ are fixed.
@@
$\varepsilon\equiv\max\left(\frac{|\Xi|}{\tilde{T}},\frac{mgL}{k_{\rm B}\tilde{T}}\right),$ (2)
@@
$x_{\rm o}=x_{\rm m}\equiv\frac{0+L}{2}$ (5)
@@
$\Psi$ is non-zero only during liquid-gas coexistence. In this case, $\Psi$ is discontinuous at $g=0$ because $X<x_{\rm m}$ for $g>0$ (assuming liquid is denser) and $X>x_{\rm m}$ for $g<0$. To determine the equilibrium configuration in liquid-gas coexistence, we construct a variational function $\mathcal{F}_{g}$, encompassing the equilibrium free energy $F_{g}$. Let $\mathrm{\ell}$ and $\mathrm{u}$ denote the lower ($0<x<x_{\theta}$) and upper ($x_{\theta}<x<L$) phases, respectively, which can be either liquid or gas. Here, $x_{\theta}$ is the interface position. Then, let ${\mathcal{V}}^{\ell}$ and ${\mathcal{N}}^{\ell}$ be the volume and particle number of the lower phase. From geometrical considerations, Eq. (7) can be expressed as
@@
${\cal F}_{g}({\mathcal{N}}^{\ell},{\mathcal{V}}^{\ell};T,V,N,mgL)\equiv F(T,{\mathcal{V}}^{\ell},{\mathcal{N}}^{\ell})$
@@
$F_{g}(T,V,N,mgL)=\min_{{\mathcal{V}}^{\ell},{\mathcal{N}}^{\ell}}{\cal F}_{g}({\mathcal{N}}^{\ell},{\mathcal{V}}^{\ell};T,V,N,mgL).$ (11)
Minimizing ${\cal F}_{g}$ with respect to $({\mathcal{N}}^{\ell},{\mathcal{V}}^{\ell})$ yields the equilibrium values $(N^{\rm L}_{\rm eq},V^{\rm L}_{\rm eq})$, corresponding to the configuration where the liquid phase occupies the lower region (assuming liquid is denser). The indices $\mathrm{L}$ and $\mathrm{G}$ specify liquid and gas, respectively.
Let $N^{\rm L}_{0}$ and $V^{\rm L}_{0}$ be the liquid amounts at $g=0$ determined by the lever rule. The saturated specific volumes $v_{\mathrm{c}}^{\mathrm{L}}(T)$ and $v_{\mathrm{c}}^{\mathrm{G}}(T)$ for liquid and gas are given by $v_{\mathrm{c}}^{\mathrm{L}}(T)=V^{\rm L}_{0}/N^{\rm L}_{0}$ and $v_{\mathrm{c}}^{\mathrm{G}}(T)=V^{\rm G}_{0}/N^{\rm G}_{0}$ with $V^{\rm G}_{0}=V-V^{\rm L}_{0}$ and $N^{\rm G}_{0}=N-N^{\rm L}_{0}$. We then have $N^{\rm L}_{\rm eq}=N^{\rm L}_{0}+O(\varepsilon)$ and $V^{\rm L}_{\rm eq}=V^{\rm L}_{0}+O(\varepsilon)$. Combining this result with the force balance equation yields the standard equilibrium phase coexistence conditions: continuity of local pressure and local chemical potential across the interface.
@@
${\cal F}_{g}({\mathcal{N}}^{\ell},{\mathcal{V}}^{\ell};\tilde{T},V,N,mgL,\Xi)\equiv F^{\mathrm{\ell}}(\tilde{T}^{\mathrm{\ell}},{\mathcal{V}}^{\ell},{\mathcal{N}}^{\ell})$
@@
with $\Psi({\mathcal{N}}^{\ell},{\mathcal{V}}^{\ell})$ given by (9). Here, $F^{\mathrm{\ell}}$ and $F^{\mathrm{u}}$ represent the free energies associated with the lower ($\mathrm{\ell}$) and upper ($\mathrm{u}$) phases depending on the global temperatures $\tilde{T}^{\mathrm{\ell}}$ and $\tilde{T}^{\mathrm{u}}$ of the respective regions
[…]
with the local free energy density $f(T(x),\rho(x))$. By definition, the global temperatures satisfy $N\tilde{T}={\mathcal{N}}^{\ell}\tilde{T}^{\mathrm{\ell}}+(N-{\mathcal{N}}^{\ell})\tilde{T}^{\mathrm{u}}$. Assuming approximately linear temperature profiles within each phase, we have $\tilde{T}^{\mathrm{\ell}}=(T_{1}+\theta)/2$ and $\tilde{T}^{\mathrm{u}}=(T_{2}+\theta)/2$, where $\theta$ is the interface temperature. Combining these relations, we obtain
$\tilde{T}^{\mathrm{\ell}}=\tilde{T}-\frac{\Xi}{2}\frac{N-{\mathcal{N}}^{\ell}}{N},\quad\tilde{T}^{\mathrm{u}}=\tilde{T}+\frac{\Xi}{2}\frac{{\mathcal{N}}^{\ell}}{N}.$ (13)
The NESS is assumed to minimize the variational function. Potential steady state configurations correspond to the stationary points of ${\cal F}_{g}$, found by solving the variational equations
$\frac{\partial{\cal F}_{g}}{\partial{\mathcal{V}}^{\ell}}=0,\qquad\frac{\partial{\cal F}_{g}}{\partial{\mathcal{N}}^{\ell}}=0,$ (14)
@@
$\mu^{\mathrm{\ell}}(T_{\rm s},p_{\theta})=\mu^{\mathrm{u}}(T_{\rm s},p_{\theta}),$ (15)
$T_{\rm s}\equiv\tilde{T}+\frac{\Xi}{2}\frac{2{\mathcal{N}}^{\ell}-N}{N},$ (16)
@@
Remarkably, Eqs. (15), (16), and (17) are consistent with the results reported for the zero-gravity ($g=0$) heat conduction case in [22, 23]. Specifically, they imply that the interface temperature $\theta$ generally deviates from the equilibrium coexistence temperature $T_{\mathrm{c}}(p_{\theta})$,
$\theta=T_{\mathrm{c}}(p_{\theta})+J\left(\frac{\bar{v}}{v_{\mathrm{c}}^{\mathrm{\ell}}\kappa^{\mathrm{u}}}-\frac{\bar{v}}{v_{\mathrm{c}}^{\mathrm{u}}\kappa^{\mathrm{\ell}}}\right)\frac{x_{\theta}(L-x_{\theta})}{2L},$ (18)
where $J$ is the heat flux (defined positive from bottom to top), $\kappa^{\mathrm{\ell},\mathrm{u}}$ are the thermal conductivities of the phases, and $\bar{v}=V/N$. This consistency suggests that the stabilization of metastable states by heat flux is a robust phenomenon, persisting even under gravity. This non-trivial prediction has been numerically confirmed for order-disorder coexistence in the Hamiltonian Potts model [24] and analytically verified for liquid-gas coexistence in boundary-driven diffusive systems [25].
@@
Note that the stationarity conditions derived from (14) generally admit two solutions: one corresponding to liquid below gas, denoted $(N^{\rm L}_{\rm s},V^{\rm L}_{\rm s})$, and the other to gas below liquid, denoted $(N^{\rm G}_{\rm s},V^{\rm G}_{\rm s})$. The NESS corresponds to the configuration that yields the global minimum of the variational function ${\cal F}_{g}$ defined in (12).
@@
${\cal F}_{g}(N^{\rm L}_{\rm s},V^{\rm L}_{\rm s})=F(\tilde{T},V,N)-\Psi(N^{\rm L}_{0},V^{\rm L}_{0})mg_{\rm eff}L,$ (19)
${\cal F}_{g}(N^{\rm G}_{\rm s},V^{\rm G}_{\rm s})=F(\tilde{T},V,N)+\Psi(N^{\rm L}_{0},V^{\rm L}_{0})mg_{\rm eff}L,$ (20)
@@
$g_{\rm eff}\equiv g+\frac{\bar{v}}{m}\frac{dp_{\mathrm{s}}}{dT}\bigg{|}_{\tilde{T}}\frac{\Xi}{L}.$ (21)
In (20), $\Psi(N^{\mathrm{G}}_{0},V^{\mathrm{G}}_{0})=-\Psi(N^{\rm L}_{0},V^{\rm L}_{0})$ was used, where $N^{\rm L}_{0}$ and $V^{\rm L}_{0}$ are the zero-gravity equilibrium amounts.
@@
${\cal F}_{g}(N^{\rm G}_{\rm s},V^{\rm G}_{\rm s})-{\cal F}_{g}(N^{\rm L}_{\rm s},V^{\rm L}_{\rm s})=2mg_{\rm eff}L\Psi(N^{\rm L}_{0},V^{\rm L}_{0}),$ (22)
is generally non-zero. The lever rule yields the expressions of $N^{\rm L}_{0}$ and $V^{\rm L}_{0}$ with the saturated specific volumes. Using them, (9) gives
$\Psi(N^{\rm L}_{0},V^{\rm L}_{0})=\frac{N}{2}\frac{(v^{\mathrm{G}}_{\mathrm{c}}(\tilde{T})-\bar{v})(\bar{v}-v^{\mathrm{L}}_{\mathrm{c}}(\tilde{T}))}{\bar{v}(v^{\mathrm{G}}_{\mathrm{c}}(\tilde{T})-v^{\mathrm{L}}_{\mathrm{c}}(\tilde{T}))}.$ (23)
For liquid-gas coexistence, we assume $v_{\mathrm{c}}^{\mathrm{L}}(\tilde{T})<\bar{v}<v_{\mathrm{c}}^{\mathrm{G}}(\tilde{T})$, which ensures $\Psi(N^{\rm L}_{0},V^{\rm L}_{0})>0$. We conclude that
${\cal F}_{g}(N^{\rm L}_{\rm s},V^{\rm L}_{\rm s})<{\cal F}_{g}(N^{\rm G}_{\rm s},V^{\rm G}_{\rm s})\qquad(g_{\mathrm{eff}}>0),$ (24)
${\cal F}_{g}(N^{\rm G}_{\rm s},V^{\rm G}_{\rm s})<{\cal F}_{g}(N^{\rm L}_{\rm s},V^{\rm L}_{\rm s})\qquad(g_{\mathrm{eff}}<0).$ (25)
The configuration with the smaller value of ${\cal F}_{g}$ corresponds to the stable NESS. The sign of $g_{\mathrm{eff}}$ dictates the configuration: liquid is at the bottom if $g_{\mathrm{eff}}>0$ and at the top if $g_{\mathrm{eff}}<0$, directly answering the question posed in the Introduction. This can be explicitly expressed by
$X=x_{\rm m}-L\frac{|g_{\mathrm{eff}}|}{g_{\mathrm{eff}}}\frac{\Psi(N^{\mathrm{L}}_{0},V^{\mathrm{L}}_{0})}{N}+O(\varepsilon)$ (26)
with (23). Figure 2(a) demonstrates how the heat flux alters the steady state under a fixed gravity. When the top is cooled sufficiently such that the mean temperature gradient $\Xi/L$ satisfies the condition $g_{\rm eff}<0$, the liquid floats above the gas. Neglecting $O(\varepsilon^{2})$, the transition point $g_{\mathrm{eff}}=0$ corresponds to the condition
@@
where $p_{1}$ and $p_{2}$ are the pressures at the bottom and the top ($p_{1}-p_{2}=mgN/A$). The criterion (27) explicitly expresses the balance between thermal driving (related to saturation pressure difference) and mechanical driving (gravity). See Appendix C for an example of experimental setup.
Equation (18) implies local states near the interface are generally metastable (i.e., $\theta\neq T_{\mathrm{c}}(p_{\theta})$). When $g>0$ and $\Xi>0$ ($T_{2}>T_{1}$), gravity and heat flux are aligned, and the gas settles at the top as expected. However, unexpectedly, a significant portion of the gas can become supercooled by the heat flux [Fig. 2(b)-(3)]. If $T_{2}$ is reduced slightly below $T_{1}$ while $g_{\mathrm{eff}}$ remains positive, the “normal” configuration (liquid at bottom) persists. If $\Xi<0$ ($T_{1}>T_{2}$) with $g_{\mathrm{eff}}>0$, the liquid-at-bottom configuration is unchanged, but phase states are altered: the gas becomes stable, and the liquid near the interface superheats [Fig. 2(b)-(2)]. This superheated liquid, absent at $g=0$, is thus a gravity-induced non-equilibrium effect. When $T_{2}$ is low enough for $g_{\mathrm{eff}}<0$, the liquid floats above the gas. Here, the liquid is stable, while the gas near the interface is supercooled [Fig. 2(b)-(1)]. While configurations (1) and (3) may appear as spatial inversions, the intermediate state (2) with superheated liquid implies a more complex interplay. Further investigation is warranted [27].
@@
$F_{g}(\tilde{T},V,N,mgL,\Xi)=\min_{{\mathcal{V}}^{\ell},{\mathcal{N}}^{\ell}}{\cal F}_{g}({\mathcal{V}}^{\ell},{\mathcal{N}}^{\ell};\tilde{T},V,N,mgL,\Xi).$ (28)
@@
We have extended global thermodynamics to unify gravity and heat flow effects into a single variational principle. Our central finding is the emergence of an effective gravity $g_{\text{eff}}$, which encapsulates their combined influence and predicts the stable phase configuration. This work also recovers the stabilization of locally metastable states near the interface, a prediction from earlier global thermodynamics studies. Deriving the thermodynamic potential $F_{g}$ from microscopic models remains a challenging task. We expect that the analytical method proposed for boundary-driven diffusive systems may be extended to address our system, and its potential application is currently under investigation. We conclude with the following remarks.
Performing molecular dynamics (MD) simulations would be valuable for quantitatively testing our predictions. We note that the phenomenon of liquid floating above gas has previously been observed numerically [28, 29]. These observed phenomena – heat-flux-induced floating of liquid against gravity – are qualitatively consistent with our predictions. The numerical data appeared to agree with the local equilibrium hypothesis at the interface. We consider that these earlier findings do not preclude the predictions derived from our variational principle, particularly regarding the deviations from local equilibrium at the interface. In fact, the MD simulations exhibited convergence that was too slow with increasing system size towards the thermodynamic limit expected from the local equilibrium hypothesis [28]. Further numerical investigations are crucial for a complete understanding of heat-conducting fluids under weak gravity.
Experimental studies on liquid-gas coexistence in heat conduction remain crucial but challenging. To our knowledge, quantitative reports on the thermodynamic properties are lacking even for the regime $\Xi>0$, where the liquid is expected at the bottom. A primary difficulty is the occurrence of unwanted convection induced by a slight temperature difference of $1$ K between the center and the periphery [30]. Precise experiments aiming to verify global thermodynamics might require careful control, such as active control [31]. Focusing on the steady interface properties or the configurational transition at $g_{\mathrm{eff}}=0$ could provide invaluable tests.
@@
The authors thank A. Yoshida, F. Kagawa, R. Higa, and A. Hisada for useful discussions. This study was supported by JSPS KAKENHI Grant Numbers JP23K22415, JP25K00923, and JP25H01975.

### v2

@@ Abstract
While a deterministic hydrodynamic description is a standard approach [8], its application to phase-coexisting systems is complicated by the singularly thin nature of the interface. This can lead to the existence of multiple stationary solutions, trapping the deterministic time evolution in a metastable state. The hydrodynamic equations alone thus fail to provide a selection principle for the physically realized NESS. Stochastic properties, as discussed in modern frameworks for NESS such as large deviation theory [9, 10] and stochastic thermodynamics [11, 12], are therefore essential for the selection of the unique steady state [13].
In equilibrium, the selection problem is resolved by a thermodynamic variational principle, such as minimizing free energy, which uniquely determines the stable state [14, 15, 16]. This motivates extending such a variational approach to non-equilibrium systems. As a variational principle in weak NESS, the minimum entropy production principle, originating from linear irreversible thermodynamics [17, 18], was proposed for the selection of a steady current [19]. Despite the mathematically precise formulation of the minimum entropy production principle [20, 21], it is still challenging to predict the true steady state among multiple steady states.
@@
We consider $N$ particles of mass $m$ confined in a rectangular container of volume $V$, height $L$, and cross-sectional area $A=V/L$. See Fig. 1. The vertical coordinate $x$ runs from the bottom wall ($x=0$) to the top wall ($x=L$). Uniform gravity $g$ acts in the negative $x$ direction, resulting in a gravitational potential energy $mg(x-x_{\mathrm{o}})$ relative to an arbitrary reference point $x_{\mathrm{o}}$. The walls at 0 and $L$ are maintained at constant temperatures $T_{1}$ and $T_{2}$, respectively. We define the temperature difference parameter $\Xi\equiv T_{2}-T_{1}$. The total particle number $N$ and volume $V$ are fixed.
@@
$x_{\mathrm{o}}=x_{\mathrm{m}}\equiv\frac{0+L}{2}$ (5)
@@
$\Psi$ is non-zero only during liquid-gas coexistence. In this case, $\Psi$ is discontinuous at $g=0$ because $X<x_{\mathrm{m}}$ for $g>0$ (assuming liquid is denser) and $X>x_{\mathrm{m}}$ for $g<0$. To determine the equilibrium configuration in liquid-gas coexistence, we construct a variational function $\mathcal{F}_{g}$, encompassing the equilibrium free energy $F_{g}$. Let $\ell$ and $\mathrm{u}$ denote the lower ($0<x<x_{\theta}$) and upper ($x_{\theta}<x<L$) phases, respectively, which can be either liquid or gas. Here, $x_{\theta}$ is the interface position. Then, let $\mathcal{V}^{\ell}$ and $\mathcal{N}^{\ell}$ be the volume and particle number of the lower phase. From geometrical considerations, (7) can be expressed as
@@
$\mathcal{F}_{g}(\mathcal{N}^{\ell},\mathcal{V}^{\ell};T,V,N,mgL)\equiv F(T,\mathcal{V}^{\ell},\mathcal{N}^{\ell})$
@@
$F_{g}(T,V,N,mgL)=\min_{\mathcal{V}^{\ell},\mathcal{N}^{\ell}}\mathcal{F}_{g}(\mathcal{N}^{\ell},\mathcal{V}^{\ell};T,V,N,mgL).$ (11)
Minimizing $\mathcal{F}_{g}$ with respect to $(\mathcal{N}^{\ell},\mathcal{V}^{\ell})$ yields the equilibrium values $(N^{\mathrm{L}}_{\mathrm{eq}},V^{\mathrm{L}}_{\mathrm{eq}})$, corresponding to the configuration where the liquid phase occupies the lower region (assuming liquid is denser). The indices $\mathrm{L}$ and $\mathrm{G}$ specify liquid and gas, respectively.
Let $N^{\mathrm{L}}_{0}$ and $V^{\mathrm{L}}_{0}$ be the liquid amounts at $g=0$ determined by the lever rule. The saturated specific volumes $v_{\mathrm{c}}^{\mathrm{L}}(T)$ and $v_{\mathrm{c}}^{\mathrm{G}}(T)$ for liquid and gas are given by $v_{\mathrm{c}}^{\mathrm{L}}(T)=V^{\mathrm{L}}_{0}/N^{\mathrm{L}}_{0}$ and $v_{\mathrm{c}}^{\mathrm{G}}(T)=V^{\mathrm{G}}_{0}/N^{\mathrm{G}}_{0}$ with $V^{\mathrm{G}}_{0}=V-V^{\mathrm{L}}_{0}$ and $N^{\mathrm{G}}_{0}=N-N^{\mathrm{L}}_{0}$. We then have $N^{\mathrm{L}}_{\mathrm{eq}}=N^{\mathrm{L}}_{0}+O(\varepsilon)$ and $V^{\mathrm{L}}_{\mathrm{eq}}=V^{\mathrm{L}}_{0}+O(\varepsilon)$. Combining this result with the force balance equation yields the standard equilibrium phase coexistence conditions: continuity of local pressure and local chemical potential across the interface.
@@
$\mathcal{F}_{g}(\mathcal{N}^{\ell},\mathcal{V}^{\ell};\tilde{T},V,N,mgL,\Xi)\equiv F^{\ell}(\tilde{T}^{\ell},\mathcal{V}^{\ell},\mathcal{N}^{\ell})$
@@
with $\Psi(\mathcal{N}^{\ell},\mathcal{V}^{\ell})$ given by (9). Here, $F^{\ell}$ and $F^{\mathrm{u}}$ represent the free energies associated with the lower ($\ell$) and upper ($\mathrm{u}$) phases depending on the global temperatures $\tilde{T}^{\ell}$ and $\tilde{T}^{\mathrm{u}}$ of the respective regions
$F^{\ell}(\tilde{T}^{\ell},\mathcal{V}^{\ell},\mathcal{N}^{\ell})=A\int_{0}^{x_{\theta}}f(T(x),\rho(x))dx,$
$F^{\mathrm{u}}(\tilde{T}^{\mathrm{u}},V-V^{\ell},N-N^{\ell})=A\int_{x_{\theta}}^{L}f(T(x),\rho(x))dx,$
with the local free energy density $f(T(x),\rho(x))$ allowing for locally metastable states. By definition, the global temperatures satisfy $N\tilde{T}=\mathcal{N}^{\ell}\tilde{T}^{\ell}+(N-\mathcal{N}^{\ell})\tilde{T}^{\mathrm{u}}$. Assuming approximately linear temperature profiles within each phase, we have $\tilde{T}^{\ell}=(T_{1}+\theta)/2$ and $\tilde{T}^{\mathrm{u}}=(T_{2}+\theta)/2$, where $\theta$ is the interface temperature. Combining these relations, we obtain
$\tilde{T}^{\ell}=\tilde{T}-\frac{\Xi}{2}\frac{N-\mathcal{N}^{\ell}}{N},\quad\tilde{T}^{\mathrm{u}}=\tilde{T}+\frac{\Xi}{2}\frac{\mathcal{N}^{\ell}}{N}.$ (13)
The NESS is assumed to minimize the variational function. Potential steady state configurations correspond to the stationary points of $\mathcal{F}_{g}$, found by solving the variational equations
$\frac{\partial\mathcal{F}_{g}}{\partial\mathcal{V}^{\ell}}=0,\qquad\frac{\partial\mathcal{F}_{g}}{\partial\mathcal{N}^{\ell}}=0,$ (14)
@@
$\mu^{\ell}(T_{\mathrm{s}},p_{\theta})=\mu^{\mathrm{u}}(T_{\mathrm{s}},p_{\theta}),$ (15)
$T_{\mathrm{s}}\equiv\tilde{T}+\frac{\Xi}{2}\frac{2\mathcal{N}^{\ell}-N}{N},$ (16)
@@
Remarkably, (15), (16), and (17) are consistent with the results reported for the zero-gravity ($g=0$) heat conduction case in [23, 24]. Specifically, they imply that the interface temperature $\theta$ generally deviates from the equilibrium coexistence temperature $T_{\mathrm{c}}(p_{\theta})$,
$\theta=T_{\mathrm{c}}(p_{\theta})+J\left(\frac{\bar{v}}{v_{\mathrm{c}}^{\ell}\kappa^{\mathrm{u}}}-\frac{\bar{v}}{v_{\mathrm{c}}^{\mathrm{u}}\kappa^{\ell}}\right)\frac{x_{\theta}(L-x_{\theta})}{2L}+O(\varepsilon^{2}),$ (18)
as a physical effect of the linear response regime $O(\varepsilon)$. Here, $J$ is the heat flux (defined positive from bottom to top), $\kappa^{\ell,\mathrm{u}}$ are the thermal conductivities of the phases, and $\bar{v}=V/N$. This consistency suggests that the stabilization of metastable states by heat flux is a robust phenomenon, persisting even under gravity. This non-trivial prediction has been numerically confirmed for order-disorder coexistence in the Hamiltonian Potts model [25] and analytically verified for liquid-gas coexistence in boundary-driven diffusive systems [13].
@@
Note that the stationarity conditions derived from (14) generally admit two solutions: one corresponding to liquid below gas, denoted $(N^{\mathrm{L}}_{\mathrm{s}},V^{\mathrm{L}}_{\mathrm{s}})$, and the other to gas below liquid, denoted $(N^{\mathrm{G}}_{\mathrm{s}},V^{\mathrm{G}}_{\mathrm{s}})$. The NESS corresponds to the configuration that yields the global minimum of the variational function $\mathcal{F}_{g}$ defined in (12).
@@
$\mathcal{F}_{g}(N^{\mathrm{L}}_{\mathrm{s}},V^{\mathrm{L}}_{\mathrm{s}})=F(\tilde{T},V,N)-\Psi(N^{\mathrm{L}}_{0},V^{\mathrm{L}}_{0})mg_{\mathrm{eff}}L,$ (19)
$\mathcal{F}_{g}(N^{\mathrm{G}}_{\mathrm{s}},V^{\mathrm{G}}_{\mathrm{s}})=F(\tilde{T},V,N)+\Psi(N^{\mathrm{L}}_{0},V^{\mathrm{L}}_{0})mg_{\mathrm{eff}}L,$ (20)
@@
$g_{\mathrm{eff}}\equiv g+\frac{\bar{v}}{m}\frac{dp_{\mathrm{s}}}{dT}\bigg|_{\tilde{T}}\frac{\Xi}{L}$ (21)
with the saturation pressure $p_{\mathrm{s}}(T)$ at temperature $T$. In (20), $\Psi(N^{\mathrm{G}}_{0},V^{\mathrm{G}}_{0})=-\Psi(N^{\mathrm{L}}_{0},V^{\mathrm{L}}_{0})$ was used, where $N^{\mathrm{L}}_{0}$ and $V^{\mathrm{L}}_{0}$ are the zero-gravity equilibrium amounts.
@@
$\mathcal{F}_{g}(N^{\mathrm{G}}_{\mathrm{s}},V^{\mathrm{G}}_{\mathrm{s}})-\mathcal{F}_{g}(N^{\mathrm{L}}_{\mathrm{s}},V^{\mathrm{L}}_{\mathrm{s}})=2mg_{\mathrm{eff}}L\Psi(N^{\mathrm{L}}_{0},V^{\mathrm{L}}_{0}),$ (22)
is generally non-zero. The lever rule yields the expressions of $N^{\mathrm{L}}_{0}$ and $V^{\mathrm{L}}_{0}$ with the saturated specific volumes. Using them, (9) gives
$\Psi(N^{\mathrm{L}}_{0},V^{\mathrm{L}}_{0})=\frac{N}{2}\frac{(v^{\mathrm{G}}_{\mathrm{c}}(\tilde{T})-\bar{v})(\bar{v}-v^{\mathrm{L}}_{\mathrm{c}}(\tilde{T}))}{\bar{v}(v^{\mathrm{G}}_{\mathrm{c}}(\tilde{T})-v^{\mathrm{L}}_{\mathrm{c}}(\tilde{T}))}$. (23)
[…]
$\mathcal{F}_{g}(N^{\mathrm{L}}_{\mathrm{s}},V^{\mathrm{L}}_{\mathrm{s}})<\mathcal{F}_{g}(N^{\mathrm{G}}_{\mathrm{s}},V^{\mathrm{G}}_{\mathrm{s}})\qquad(g_{\mathrm{eff}}>0),$ (24)
$\mathcal{F}_{g}(N^{\mathrm{G}}_{\mathrm{s}},V^{\mathrm{G}}_{\mathrm{s}})<\mathcal{F}_{g}(N^{\mathrm{L}}_{\mathrm{s}},V^{\mathrm{L}}_{\mathrm{s}})\qquad(g_{\mathrm{eff}}<0).$ (25)
The configuration with the smaller value of $\mathcal{F}_{g}$ corresponds to the stable NESS. The sign of $g_{\mathrm{eff}}$ dictates the configuration: liquid is at the bottom if $g_{\mathrm{eff}}>0$ and at the top if $g_{\mathrm{eff}}<0$, directly answering the question posed in the Introduction. This can be explicitly expressed by
$X=x_{\mathrm{m}}-L\frac{|g_{\mathrm{eff}}|}{g_{\mathrm{eff}}}\frac{\Psi(N^{\mathrm{L}}_{0},V^{\mathrm{L}}_{0})}{N}+O(\varepsilon)$ (26)
with (23). Figure 2(a) demonstrates how the heat flux alters the steady state under a fixed gravity. This stabilization occurs when the entropic contribution to the variational function $\mathcal{F}_{g}$, quantified by the thermal term in $g_{\mathrm{eff}}$, overcomes the increase in gravitational potential energy.
When the top is cooled sufficiently such that the mean temperature gradient $\Xi/L$ satisfies the condition $g_{\mathrm{eff}}<0$, the liquid floats above the gas. The transition point $g_{\mathrm{eff}}=0$ corresponds to the condition
@@
where $p_{1}$ and $p_{2}$ are the pressures at the bottom and the top. This follows from (21) with the relations $(dp_{\mathrm{s}}/dT)\Xi=p_{\mathrm{s}}(T_{2})-p_{\mathrm{s}}(T_{1})+O(\varepsilon^{2})$ and $p_{1}-p_{2}=mgL/\bar{v}$. The criterion (27) explicitly expresses the balance between thermal driving (related to saturation pressure difference) and mechanical driving (gravity). See Appendix C for an example of experimental setup.
When $g>0$ and $\Xi>0$ ($T_{2}>T_{1}$), gravity and heat flux are aligned, and the gas settles at the top as expected. However, unexpectedly, a significant portion of the gas can become supercooled by the heat flux as shown in (18) [Fig. 2(b)-(3)]. If $T_{2}$ is reduced slightly below $T_{1}$ while $g_{\mathrm{eff}}$ remains positive, the “normal” configuration (liquid at bottom) persists. If $\Xi<0$ ($T_{1}>T_{2}$) with $g_{\mathrm{eff}}>0$, the liquid-at-bottom configuration is unchanged, but phase states are altered: the gas becomes stable, and the liquid near the interface superheats [Fig. 2(b)-(2)]. This superheated liquid, absent at $g=0$, is thus a gravity-induced non-equilibrium effect. When $T_{2}$ is low enough for $g_{\mathrm{eff}}<0$, the liquid floats above the gas. Here, the liquid is stable, while the gas near the interface is supercooled [Fig. 2(b)-(1)]. While configurations (1) and (3) may appear as spatial inversions, the intermediate state (2) with superheated liquid implies a more complex interplay. Further investigation is warranted [27].
@@
$F_{g}(\tilde{T},V,N,mgL,\Xi)=\min_{\mathcal{V}^{\ell},\mathcal{N}^{\ell}}\mathcal{F}_{g}(\mathcal{V}^{\ell},\mathcal{N}^{\ell};\tilde{T},V,N,mgL,\Xi).$ (28)
@@
We have extended global thermodynamics to unify gravity and heat flow, yielding a variational principle governed by an effective gravity, $g_{\mathrm{eff}}$, that predicts the stable phase configuration. This work also confirms the stabilization of local metastable states, a prediction of earlier global thermodynamics studies. Deriving our thermodynamic potential from microscopic models remains an important future challenge.
Our predictions could be tested by molecular dynamics (MD) simulations. Indeed, the liquid floating phenomenon has been observed numerically [28, 29], in qualitative agreement with our theory. While these studies assumed local equilibrium at the interface, their slow convergence towards the thermodynamic limit suggests that deviations from local equilibrium, as predicted by our principle, may play a crucial role.
Experimental verification is crucial but challenging. Quantitative reports are lacking, primarily due to unwanted convection, which can be induced by horizontal temperature differences as small as $1$ K [30]. Precise experiments to test global thermodynamics would require careful control, such as active control [31], and should focus on the steady interface properties or the configurational transition at $g_{\mathrm{eff}}=0$.
@@
The authors thank A. Yoshida, F. Kagawa, R. Higa, and A. Hisada for useful discussions. This study was supported by JSPS KAKENHI Grant Numbers JP23K22415, JP25K00923, JP25K22002, and JP25H01975.